OpenStudy (anonymous):

Solve the equation. 4(x - 1.6) = 9.3

OpenStudy (anonymous):

4(x - 1.6) = 9.3 (x - 1.6) = 9.3/4 x - 1.6 = 2.325 x = 3.925